# Knowledge-guided machine learning model with soil moisture for corn yield prediction under drought conditions

### v1

@@ 5.1 Spatial influence of soil moisture on model prediction
We visualized the attention scores of soil moisture across twelve states in the U.S. Corn Belt from June to August over the years 2019 to 2023 (Fig. 8). The attention scores indicated the relative importance assigned to soil moisture by the model in different regions, with higher scores suggesting a stronger influence on yield prediction. The attention map highlighted how the model’s reliance on soil moisture varied across different growth stages and drought conditions, allowing us to assess whether soil moisture has a greater impact on the model in drought-affected areas.
@@
In June, during the early growth period, the attention scores for soil moisture were generally lower across all four years, suggesting that soil moisture is less critical at this stage. This might be due to cooler temperatures and lower evaporation rates, which reduced the immediate need for water. In July, attention scores increased, particularly in drought-affected areas, as soil moisture became more influential during active growth and vegetative stages. By August, attention scores peaked in drought-affected states, aligning with the critical reproductive phase of corn when adequate soil moisture was essential for kernel development.
From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations, aligning with the drought severity in different years. Note that, for better visualization, we normalized the attention values within each year. Therefore, we only need to focus on the differences across regions within the same year, rather than making cross-year comparisons. In early 2019 (Fig. 8(a1)), the attention was concentrated in Iowa, reflecting the drought-affected states. Subsequently (Fig. 8(a3)), the regions with higher attention values gradually expanded, but Iowa remained the area with the highest attention. In 2020 (Fig. 8(b)), the high-attention regions expanded to include North Dakota, South Dakota, and Nebraska, corresponding to widespread drought in USDA drought map. In 2021 (Fig. 8(c)), the attention intensified over North Dakota, South Dakota, Minnesota, and Nebraska, mirroring severe drought conditions. By 2022 (Fig. 8(d)), high-attention areas were mainly in Nebraska, Kansas, South Dakota, and Iowa. Additionally, over time, there was a trend of drought expanding toward the central region. In 2023 (Fig. 8(e)), the attention peaked across Nebraska, Kansas, Minnesota, Iowa, Wisconsin, and Missouri. Finally, in August, it reached its peak, and soil moisture became more important in most states in the central region.
This above analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions. The elevated attention scores in drought-prone regions suggested that soil moisture serves as a crucial predictor for accurately estimating corn yield in water-stressed environments.
@@ 5.2 Temporal role of soil moisture during the corn growth season
To illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—every 16 days from June to September. We found that the VIs had the highest influence on the model around August, which aligned with findings from previous studies (Johnson, 2014). In this period, VIs shows the strongest correlation with corn yield because this period corresponds to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield. High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation. Additionally, we found that weather data showed a significant increase in attention around July in 2020 and 2022, and the periods with noticeable attention spikes aligned with the trends of VIs. This phenomenon was more pronounced in soil moisture, as soil moisture attention showed a strong correlation with VIs in all years except 2022. In 2022, a slight increase in soil moisture attention could still be observed near the VIs peak. This may be because VIs during these periods are closely related to certain weather data and soil moisture data. For example, NDWI increases with higher PPT because more rainfall enhances soil moisture and plant water content, GCI decreases with excessive Tmax because extreme heat can degrade chlorophyll and reduce plant vitality, etc. This indicated that our attention mechanism effectively captured feature importance over time dimension.
@@ 5.3 Statistical impact of soil moisture in drought and non-drought regions
The box plot illustrates the distribution of soil moisture attention across all counties from 2019 to 2023 (Fig. 10). Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought regions and blue-marked non-drought regions from 2019 to 2023. Through this visualization, we aimed to analyze the differences in soil moisture attention between drought-affected and non-drought counties across all years. In 2021, the most severe drought year, both the variance and magnitude of soil moisture attention were the highest. We also observed that attention values were notably high in 2019, which aligns with our experimental results, where our KGML-SM model achieved the most significant improvement over other models. Since our model better captured the impact of soil moisture, it achieved the most significant improvement in years when soil moisture has a greater influence. In contrast, the values for 2020, 2022, and 2023 were more concentrated, with fewer extreme values. Overall, in all years, soil moisture attention in drought-affected regions was consistently higher than in non-drought regions, indicating that soil moisture had a greater impact on the model in these areas.
@@ 6 Conclusion
In this study, we proposed the KGML-SM model, which first supplements the limited soil moisture data with simulated data. It then learned the PB relationship between weather and soil moisture. Additionally, based on prior knowledge that drought leads to yield overestimation, we introduced a drought-aware loss function to mitigate this issue in drought-affected areas. We analyzed 12 states in the U.S. Corn Belt to examine the impact of soil moisture on corn yield prediction. Our method consistently outperformed baseline models across multiple test years. We also explore the spatial and temporal impact of soil moisture on the model by visualizing attention, highlighting when and where the model focuses more on soil moisture. In the future, we aim to apply transfer learning to adapt models trained on well-studied regions with simulated data to areas with limited data availability.

### v2

@@ 5.1 Spatial influence of soil moisture on model prediction
To answer the first question, we visualized the attention scores of soil moisture across twelve states in the U.S. Corn Belt from June to August over the years 2019 to 2023 (Fig. 8). The attention scores indicated the relative importance assigned to soil moisture by the model in different regions, with higher scores suggesting a stronger influence on yield prediction. The attention map highlighted how the model’s reliance on soil moisture varied across different growth stages and drought conditions, allowing us to assess whether soil moisture has a greater impact on the model in drought-affected areas. To enhance visualization, we normalized attention values within each year. Consequently, analysis should focus on attention trends across regions within the same year rather than comparing values across different years.
@@
In June, during the early growth period, the attention scores for soil moisture were generally lower across all five years, This might be because the temperature was not too high and drought conditions were not severe at this time, leading to a less pronounced correlation between soil moisture and corn yield. In July, attention scores increased, particularly in drought-affected areas, as soil moisture became more influential during active growth and vegetative stages. By August, attention scores peaked in drought-affected states, aligning with the critical reproductive phase of corn when adequate soil moisture was essential for kernel development.
From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations, aligning with the drought severity in different years. In early 2019 (Fig. 8(a1)), attention concentrated in Iowa. Subsequently (Fig. 8(a3)), regions with higher attention values expanded gradually, though Iowa retained the highest attention, aligning with observed drought conditions that year. During 2020 (Fig. 8(b)), high-attention regions were initially concentrated in Iowa and Nebraska, later expanding to adjacent areas. By 2021 (Fig. 8(c)), attention intensified over North Dakota, South Dakota, Minnesota, and Nebraska. In 2022 (Fig. 8(d)), high-attention areas were primarily located in Nebraska, Kansas, South Dakota, and Iowa. In 2023 (Fig. 8(e1)), attention initially focused on Kansas, expanding thereafter across a majority of states. Ultimately, significant attention emerged in Nebraska, Kansas, Minnesota, Iowa, Wisconsin, and Missouri, indicative of substantial drought conditions (Fig. 8(e3)). Overall, the period from 2019 to 2023 exhibited a trend of drought conditions and corresponding attention expanding towards the central United States, consistent with actual observations. The reason for the incomplete consistency between the two could be that although drought was present in the region, it did not severely impact corn yield. Alternatively, even without drought in a particular area, other factors, such as the derecho storm, might have affected corn yield. This analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions.
@@ 5.2 Temporal role of soil moisture during the corn growth season
To answer the second question and illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—every 16 days from June to September in drought-affected states (Fig. 9). We found that the VIs had the highest influence on the model around August, which aligned with findings from previous study (Johnson, 2014). In this period, VIs showed the strongest correlation with corn yield because this period corresponded to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield (Huang et al., 2014). High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation (Wang et al., 2005). Additionally, we found that weather data showed a significant increase in attention around July in 2020 (Fig. 9(b)) and 2022 (Fig. 9(d)), and the periods with noticeable attention spikes aligned with the trends of VIs. This phenomenon was more pronounced in soil moisture, as soil moisture attention showed a strong correlation with VIs in all years except 2022. In 2022, a slight increase in soil moisture attention could still be observed near the VIs peak. This might be because VIs during these periods are closely related to certain weather data and soil moisture data. For example, NDWI increases with higher PPT because more rainfall enhances soil moisture and plant water content (Ashok et al., 2021). This indicated that our attention mechanism effectively captured feature importance over time dimension.
@@ 5.3 Statistical impact of soil moisture in drought and non-drought regions
To answer the third question, The box plot was used to illustrate the distribution of soil moisture attention across all counties from 2019 to 2023 (Fig. 10). Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought-affected areas and blue-marked non-drought areas from 2019 to 2023 based on Table 7. Across all five years of attention values, drought-affected areas exhibited fewer outliers compared to non-drought areas, suggesting that the model’s attention to soil moisture is more stable in drought-affected regions. Additionally, in all years, the median of soil moisture attention in drought-affected areas was consistently higher than in non-drought areas, indicating that soil moisture had a greater impact on the model in these areas.
@@
5.4 Interpreting corn yield prediction errors via soil moisture prediction anomalies
To answer the fourth question, we plotted the absolute error map of both soil moisture and corn yield predictions (Fig. 11). In our modeling process, we treated soil moisture as an intermediate variable to guide the corn yield output, so when corn yield prediction was inaccurate, the corresponding soil moisture might also exhibit anomalies. This analysis aims to conduct interpretability analysis of inaccurate corn yield predictions. In 2019 (Fig. 11(a)), the drought-affected state Michigan exhibited relatively accurate corn yield predictions from the KGML-SM model; however, the soil moisture output was less accurate. This suggested that soil moisture was challenging to predict in drought-affected areas. In 2020 (Fig. 11(b)), inaccurate corn yield prediction in Iowa was primarily attributable to the derecho storm’s impact (Hosseini et al., 2020), which soil moisture could not reflect. In 2021 (Fig. 11(c)), drought in North and South Dakota resulted in inaccurate corn yield predictions, and corresponding soil moisture predictions were also found to be inaccurate. In 2022 (Fig. 11(d)), Nebraska, Kansas, and South Dakota experienced drought, and their soil moisture predictions were also inaccurate. In 2023 (Fig. 11(e)), Iowa and Missouri were affected by drought. Iowa’s corn yield and soil moisture prediction were accurate, which indicated that our model performed well. But the prediction for Missouri exhibited errors, despite relatively accurate corresponding soil moisture predictions for the state. According to report (Ward, 2023), some damaging winds and hail significantly impacted corn field, causing root lodging in late June and early July, particularly around Mooresville and Rockport, Missouri. Hail damage was also reported at Garden City, Missouri. These types of mechanical damage were not reflected in standard soil moisture data and constituted yield-limiting factors likely contributing to the corn yield prediction inaccuracies in Missouri. The analysis indicates that a significant portion of the error in corn yield prediction is related to soil moisture prediction error, given the key role soil moisture plays in corn growth; thus, if soil moisture is not accurately predicted, the accumulated error in modeling will lead to inaccurate corn yield prediction.
Figure 11: The absolute error map for (1) corn yield prediction and (2) soil moisture prediction from (a) 2019 to (e) 2023.
@@ 6 Conclusion
In this study, we proposed the KGML-SM model, where the W2S encoder is designed to capture the influence of weather on soil moisture, and the attention module is employed to weight different input features for final corn yield prediction. To address the known issue that drought conditions often lead to yield overestimation, we introduced a drought-aware loss function to mitigate this problem in drought-affected regions. We constructed both an APSIM field-level dataset and a GEE county-level dataset, leveraging a process-based model to learn the corn growth process by pretraining the KGML-SM model on the APSIM field-level dataset and finetuning it on the GEE county-level dataset. Our analysis covered 12 states in the U.S. Corn Belt to investigate the impact of soil moisture on corn yield prediction. The proposed method consistently outperformed baseline models across multiple test years. Furthermore, we studied the spatial and temporal influence of soil moisture through attention visualization, revealing when and where the model places greater focus on soil moisture. Finally, based on the relationship between soil moisture and corn yield prediction, we investigated the causes of inaccuracies in corn yield prediction and provided an explanation. In future work, we aim to apply transfer learning techniques to adapt models trained on well-studied regions with abundant simulated data to regions with limited data availability.
Funding
This work was supported by the United States Department of Agriculture (USDA) National Institute of Food and Agriculture (NIFA) Agriculture and Food Research Initiative Foundational Program (Award No. 2022-67021-36468); and the USDA NIFA Hatch Project (Accession No. 7005141).